OpenStudy (anonymous):

what does the g in the equation -4=g+1 symbolize???

OpenStudy (anonymous):

are you asking what the value is? g+1=-4 g+1-1=-4-1 g=-4-1 g=-5

OpenStudy (anonymous):

Its a symbol of the variable Whenever it change the result changes